A company has just paid a dividend of $ 3 per share, D0=$ 3 . It is estimated that the company's dividend will grow at a rate of 18 % percent per year for the next 2 years, then the dividend will grow at a constant rate of 7 % thereafter. The company's stock has a beta equal to 1.4, the risk-free rate is 4.5 percent, and the market risk premium is 4 percent. What is your estimate of the stock's current price? Round your answer to two decimal places.
Answer #1

Last Dividend, D0 = $3.00

Growth rate for next 2 years is 18% and a constant growth rate (g) of 7%

D1 = $3.00 * 1.18 = $3.54
D2 = $3.54 * 1.18 = $4.1772
D3 = $4.1772 * 1.07 = $4.469604

Required Return, rs = Risk-free Rate + Beta * Market Risk Premium
Required Return, rs = 4.50% + 1.40 * 4.00%
Required Return, rs = 10.10%

P2 = D3 / (rs - g)
P2 = $4.469604 / (0.1010 - 0.07)
P2 = $144.1808

P0 = $3.54/1.1010 + $4.1772/1.1010^2 + $144.1808/1.1010^2
P0 = $125.60

Therefore, the estimate of the stock current price is $125.60